a) For a gas reaction at 400K, the rate is reported as -dPA/dt= 3.66 PA^2. What is the value of rate constant for this reaction if the rate equation is measured in terms of mol/ (l. h)?


1
Expert's answer
2021-09-07T05:56:50-0400

First, we establish that if the rate is constant, then:

-rA = -(1/V) DnA/dt = kCA , mol/liter-hr

k = 3.66 = - (dPA /dt) / pA2 = (Atm/hr)2 = (1/Atm)/ (Atm)2

We understand that:

pA=CA  R T or dPA = (RT) dCA

With the provided figures, we determine the value of ‘k’ as:

- (RT) dCA/dt = 3.66 (CA R T)2

= 3.66 (0.0820575) (400) = 120.1322

        We get: 120.1322 lit/mol-hr